# Salient Object Detection From Arbitrary Modalities

### v1

@@ II-B RGB-D SOD
Recently, RGB-D SOD has witnessed remarkable advancements and lots of works have been published. According to their ways of exploiting cross-modal complementary information, most existing RGB-D SOD can be roughly categorized into three classes, i.e. pixel-level fusion, feature-level fusion, and decision-level fusion.
@@ II-C RGB-T SOD
The research on RGB-T SOD started relatively late compared to the above two ones. Until 2019, the first deep learning-based RGB-D SOD model [8] emerged. In the last two years, RGB-T SOD has attracted significant research interest. Generally speaking, most existing RGB-T SOD [33, 34, 35, 36, 37, 38, 39] models are based on feature-level fusion. Most existing models primarily focus on enhancing multi-modal feature fusion strategies and refining saliency strategies. For example, Chen et al. [35] introduced a modality transfer fusion (MTF) module, which first reduces the semantic gap between single and multi-modal images, and then captures the cross-modal complementary information by leveraging some point-to-point structural similarity information. While Wang et al. [36] proposed a Cross-Scale Alternate Guiding Fusion (CSAGF) module in the prediction stage, which effectively explores and combines high-level features from various scales.
@@ II-D Modality Unified SOD
In recent years, the use of complementary information, i.e., depth or thermal information, has shown its benefits to salient object detection (SOD). However, the RGB-D or RGB-T SOD problems are mainly solved independently. With the development of SOD, some researchers tend to simultaneously achieve the two types of multi-modality SOD tasks (i.e.,RGB-D and RGB-T SOD tasks) in a unified framework. For example, Gao et al. [40, 41] proposed the first unified end-to-end framework for both RGB-D and RGB-T SOD, where three main issues are jointly analyzed and solved, including (1) How to select and utilize available unimodal features; (2) How to combine cross-modal information more effectively and accurately; (3) How to process the situations where multiple objects have very different sizes in a scene. In some others computer vision fields, modality unified framworks have also aroused the concern of researchers and made impressive progress[42]. Besides, VSCode [43] explores the relations among SOD task and camouflaged object detection (COD) task and bulids a unfiled framework for four SOD tasks and two COD tasks, recently.
Although such modality unified frameworks show their robustness and generality on different modality combination(e.g., RGB-D and RGB-T), they haves to be retrained to meet the task requirements when the input modality combinations change. Differently, our proposed AM SOD model aims at taking the images of arbitrary modality types and arbitrary modality numbers as the inputs and robustly output the corresponding salient objects without additional training costs.
@@ IV-E3 Qualitative results for our proposed MSN with different types of inputs
TABLE VII: Quantitative results of each component of our proposed model.
@@
IV-E4 Effectiveness of dynamic fusion
@@
IV-E5 Visualization of modality indicator weights
@@ V Future works
For the AM SOD task, we propose a preliminary solution, i.e. MSN, which obtains some good results, but still faces many issues. Especially, our proposed model cannot achieve the best performance in many settings. This may be due to the following reasons. First, our feature extractor, MSFE, is based on CNN network, which extracts unimodal features from different modalities in a feature selection way. This implicitly narrows the feature space for representing different modalities. Actually, the idea of using modality prompts on top of Transformer in VScode [43] may be an better way for addressing such issues. Secondly, our training strategy is also sub-optimal, which does not allow the interactions among one-modality inputs, two-modality inputs and three-modality inputs. We also eagerly anticipate that a greater number of scholars will embark on exploring the topic and contribute to the advancement of this field. Accordingly, we will provide access to the code of our proposed MSN and our MSN datasets after the article is published. Please follow our github web page at ’https://github.com/nexiakele/AMSODFirst’. Such information has also been added to our revision.

### v2

@@ II-B RGB-D SOD
Recently, RGB-D SOD has witnessed remarkable advancements and lots of works have been published [20]. According to their ways of exploiting cross-modal complementary information, most existing RGB-D SOD can be roughly categorized into three classes, i.e. pixel-level fusion, feature-level fusion, and decision-level fusion.
@@ II-C RGB-T SOD
The research on RGB-T SOD started relatively late compared to the above two ones. Until 2019, the first deep learning-based RGB-T SOD model [8] emerged. In the last two years, RGB-T SOD has attracted significant research interest. Generally speaking, most existing RGB-T SOD [34, 35, 36, 37, 38, 39, 40] models are based on feature-level fusion. Most existing models primarily focus on enhancing multi-modal feature fusion strategies and refining saliency strategies. For example, Chen et al. [36] introduced a modality transfer fusion (MTF) module, which first reduces the semantic gap between single and multi-modal images, and then captures the cross-modal complementary information by leveraging some point-to-point structural similarity information. While Wang et al. [37] proposed a Cross-Scale Alternate Guiding Fusion (CSAGF) module in the prediction stage, which effectively explores and combines high-level features from various scales.
@@ II-D Modality Unified SOD
In recent years, the use of complementary information, i.e., depth or thermal information, has shown its benefits for salient object detection (SOD). However, the RGB-D or RGB-T SOD problems are mainly solved independently. With the development of SOD, some researchers tend to simultaneously achieve the two types of multi-modal SOD tasks (i.e., RGB-D and RGB-T SOD tasks) in a unified framework. For example, Gao et al. [41, 42] proposed the first unified end-to-end framework for both RGB-D and RGB-T SOD. Although such modality unified frameworks show their robustness and generality on different modality combinations (e.g., RGB-D and RGB-T), they have to be retrained to meet the task requirements when the input modality combinations change. Differently, our proposed AM SOD model aims to take images of arbitrary modality types and arbitrary modality numbers as inputs and robustly output the corresponding salient objects without additional training costs.
Recently, Jia et al. [43] presented an all-in-one salient object detection model, which can process RGB SOD tasks, RGB-T SOD tasks, and RGB-D SOD tasks by using one model. Besides, in some other computer vision fields, modality unified frameworks have also aroused the interest of researchers and made impressive progress [44, 45]. For example, VSCode [45] explores the relations between SOD tasks and camouflaged object detection (COD) tasks and builds a unified framework for four SOD tasks and two COD tasks, recently. However, those works mainly focus on the aspect of task unified paradigm in the output terminal, aiming at solving different tasks by using one framework. Different from them, we mainly focus on the aspect of modality unified paradigm for SOD task, which pays more attention to the input terminal. Specifically, although they may seem to be similar, we do not try to design a unified paradigm for different SOD tasks, such as RGB SOD, RGB-T SOD and RGB-D SOD. Alternatively, we try to detect salient objects from arbitrary modalities, whose task is fixed but its inputs are optional or unified for different modalities. Accordingly, this may provide more perspectives for other visual tasks.
@@ IV-E3 Qualitative results for our proposed MSN with different types of inputs
IV-E4 Effectiveness of MSFE
As shown in Table. VII, we verify the effectiveness of our proposed MSFE. ‘$w/o$ MID’ does not employ the modality indicator. ‘$w$ MID + $w/o$ MIWG’ directly generates modality indicator weights from the input modality indicators without using MIWG. ’$w$ MID + $w$ MIWG’ denotes the model that further employs MIWG on top of ‘$w$ MID + $w/o$ MIWG’. ‘$w$ MID + $w$ MIWG + $w$ $L_{modality}$’ denotes the final version of MSFE by further employing the cross-modality loss function $L_{modality}$.
The resuts of ‘$w$ MID + $w/o$ MIWG’ indicate that directly generating feature selection weights cannot effectively select discriminative unimodal features according to the input modalities, since the input modality indicators are unrelated to those extracted features. Furthermore, by virtue of MIWG, ‘$w$ MID + $w/o$ MIWG’ can explore the interactions between the modality indicators with those extracted features, thus abling to select discriminative unimodal features according to the input modalities. Finally, our proposed MSFE can generate more modality-distinguishable weights for unimodal feature selection by using the cross-modality loss function $L_{modality}$.
TABLE VII: Quantitative results from different variants of our proposed MSFE.
@@ IV-E4 Effectiveness of MSFE
IV-E5 Effectiveness of dynamic fusion
@@
IV-E6 Visualization of modality indicator weights
@@ V Future works
For the AM SOD task, we propose a preliminary solution, i.e. MSN, which obtains some good results, but still faces many issues. Especially, our proposed model cannot achieve the best performance in many settings. This may be due to the following reasons. First, our feature extractor, MSFE, is based on CNN network, which extracts unimodal features from different modalities in a feature selection way. This implicitly narrows the feature space for representing different modalities. Actually, the idea of using modality prompts on top of Transformer in VScode [45] may be an better way for addressing such issues. Secondly, our training strategy is also sub-optimal, which does not allow the interactions among one-modality inputs, two-modality inputs and three-modality inputs. We also eagerly anticipate that a greater number of scholars will embark on exploring the topic and contribute to the advancement of this field. Accordingly, we will provide access to the code of our proposed MSN and our AM-XD datasets after the article is published. Please follow our github web page at ’https://github.com/nexiakele/AMSODFirst’. Such information has also been added to our revision.